# DuSEGO: Dual Second-order Equivariant Graph Ordinary Differential Equation

### v1

@@ Abstract.
Graph Neural Networks (GNNs) with equivariant properties have achieved significant success in modeling complex dynamic systems and molecular properties. However, their expressiveness ability is limited by: (1) Existing methods often overlook the over-smoothing issue caused by traditional GNN models, as well as the gradient explosion or vanishing problems in deep GNNs. (2) Most models operate on first-order information, neglecting that the real world often consists of second-order systems, which further limits the model’s representation capabilities. To address these issues, we propose the Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO) for equivariant representation. Specifically, DuSEGO apply the dual second-order equivariant graph ordinary differential equations (Graph ODEs) on graph embeddings and node coordinates, simultaneously. Theoretically, we first prove that DuSEGO maintains the equivariant property. Furthermore, we provide theoretical insights showing that DuSEGO effectively alleviates the over-smoothing problem in both feature representation and coordinate update. Additionally, we demonstrate that the proposed DuSEGO mitigates the exploding and vanishing gradients problem, facilitating the training of deep multi-layer GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to baselines.
@@ 1. Introduction
GNNs (kipf2017semisupervised, ; Yao2023ImprovingTE, ; yin2024dynamic, ; yin2023coco, ) with equivariant properties (satorras2021n, ; han2022equivariant, ; liu2024segno, ) have become essential tools in modeling complex dynamic systems (karplus2002molecular, ; huang2021equivariant, ; liu2024dynamic, ; lu2024pascl, ) and molecular properties (brandstetter2021geometric, ; meng2024heterogeneous, ) due to their ability to capture symmetrical relationships within data. Equivariant GNNs are particularly effective in scenarios when maintaining consistency under transformations (e.g., rotations, reflections) is crucial. The consistency of equivariant representations allows these models to accurately reflect the inherent properties of the systems they are designed to study, leading to more reliable and insightful predictions, which now is an essential topic in machine learning (battaglia2016interaction, ; thomas2018tensor, ; zhang2022would, ; zhang2023investigating, ; meng2024deep, ).
Despite their successes, existing Equivariant GNNs face notable limitations that constrain their performance and expressiveness. First, the over-smoothing problem (li2018deeper, ; keriven2022not, ; min2020scattering, ; yin2023omg, ). The classical Equivariant GNNs (such as EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, )) typically utilize the GNNs (kipf2017semisupervised, ; yin2022generic, ; ai2023gcn, ; shou2023adversarial, ) as a backbone, as well as cooperating coordinates for equivariant representation. However, GNNs usually suffer from the over-smoothing problem, leading to the indistinguishable representation of features and coordinates, which is shown in Fig. 1. Second, gradient explosion or vanishing problems (dasoulas2021lipschitz, ; guo2022orthogonal, ; yinsport, ). Motivated by the notion that larger models are usually more expressive (kaplan2020scaling, ), we tend to explore the expressiveness ability by enlarging the model scale. However, as the network depth increases, gradients can either become excessively large or diminish to near zero, which limits the scalability and depth of GNNs. Third, real-world simulation (norcliffe2020second, ). Most Equivariant GNNs (fuchs2020se, ; thomas2018tensor, ; satorras2021n, ; brandstetter2021geometric, ; liu2023graph, ; liu2022mlp, ) primarily focus on first-order information. However, many physical systems are governed by second-order laws (norcliffe2020second, ). The first-order methods overlook the higher-order interactions that are critical in real-world systems, and limit the representational ability. Even though some works (huang2021equivariant, ; liu2023improving, ) try to introduce second-order information into equivariant learning, they simply focus on the simulation the position in dynamic physical system, ignoring the expressiveness representation for molecular properties prediction.
@@
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions.
Figure 1. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions $\mathbf{X}^{n}$ propagated through a EGNN (satorras2021n, ), SEGNN (brandstetter2021geometric, ) and GMN (huang2021equivariant, ) on a N-body system dataset, where we give the definition of the Dirichlet energy in Eq. 4.
To address these challenges, we take an insight into the second-order method in Equivariant GNNs and a framework named Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO). The proposed DuSEGO is a universal framework that transfer the first-order Equivariant GNNs to second-order, which aims to enhance Equivariant GNNs expressiveness and address the issues of over-smoothing and gradient stability. First, we prove that the proposed DuSEGO maintains the equivariant properties. Then, we analyze the properties of DuSEGO to reveal how the proposed method solves the over-smoothing and gradient exploding and vanishing problems, facilitating the training of deep GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to existing baselines.
The main contributions can be summarized as three-fold:
@@ 2. Related Work
Equivariant Graph Neural Networks. GNNs (kipf2017semisupervised, ; pang2023sa, ; yin2023messages, ; wang2024degree, ) with equivariant properties have demonstrated their effectiveness in various tasks, especially for 3D molecular graphs and physical symmetry of systems (zhang2023universal, ). The Tensor field networks (TFNs) (thomas2018tensor, ) method was first proposed for 3D point clouds, which only leverage the positional information of nodes without any features. SE(3)-transformers (fuchs2020se, ) deploy attention mechanism into TFNs to improve the model performance. EGNN (satorras2021n, ), a simple but effective form of equivariant graph neural network, was proposed and has been the backbone of the following research. SEGNN (brandstetter2021geometric, ) incorporates geometric and physical properties into the message-passing framework, significantly enhancing the accuracy and predictive capabilities of EGNN in understanding three-dimensional molecular structures. However, these methods focus on first-order velocity information, despite the fact that many physical systems are governed by second-order motion laws. GMN (huang2021equivariant, ) first introduced the second-order information into equivariant learning, and SEGNO (liu2023improving, ) further applied the second-order ODE for solving the physical dynamic system. However, these methods only target the description of positional state information in the physical world and cannot be applied to molecular feature representation learning. The proposed DuSEGO utilizes the second-order ODE to evaluate the evolution of equivariant representation and position, making it suitable for applications in dynamic physical systems and molecular properties prediction.
@@
Graph Ordinary Differential Equations. Graph Ordinary Differential Equations (Graph ODEs) (poli2019graph, ) draw inspiration from Neural Ordinary Differential Equations (chen2018neural, ) and provide a sophisticated mechanism for modeling dynamic changes within graph structures, showing great potential in graph learning (bhattoo2023learning, ; poli2019graph, ; sanchez2019hamiltonian, ; bishnoi2022enhancing, ). A series of works propose to deploy first-order or second-order ODE into GNNs to enhance their expressiveness ability. LG-ODE (huang2020learning, ) learns the dynamics of continuous systems by using models that can handle irregular sampling and incomplete data, improving prediction accuracy and robustness. CG-ODE (huang2021coupled, ) uses coupled ordinary differential equations on graphs to learn and model the dynamics of interacting systems effectively. HOPE (luo2023hope, ) leverages second-order differential equations to capture complex interactions and dynamic behaviors within graph structures. GraphCon (rusch2022graph, ) proposes a novel framework that integrates the dynamics of coupled oscillators with graph theory and analyzes the stability across complex networked systems. However, the above works do not consider the equivariance contained in the graph structure, which is hard to extend into 3D molecular graphs and the physical symmetry of systems.
@@ 3. Preliminary
Equivariance Let $T_{g}:X\rightarrow X$ be a set of transformations on $X$ for the abstract group $g\in G$. We say a function $\phi:X\rightarrow Y$ is equivariant to $g$ if there exists an equivalent transformation on its output space $S_{g}:Y\rightarrow Y$ such that: $\phi(T_{g}(\mathbf{x}))=S_{g}(\phi(\mathbf{x}))$. The equivariance properties include translations, rotations, and permutation, which ensures that the predictions of a model are consistent regardless of the orientation or position of the objects within the input data.
@@
Graph Nerual Networks Given graph $G=(\mathcal{V},\mathcal{E})$, let $\mathbf{h}_{v}^{(k)}$ denote embedding vector for node $v$ at the $l$-th layer. For node $v\in\mathcal{V}$, we aggregate the embeddings from the neighbors of $v$ at layer $l-1$. Then, the node embedding $\mathbf{h}_{v}^{(k)}$ is iteratively updated by combining the embedding of $v$ in the previous layer with the aggregated neighbor embedding. In formulation, $\mathbf{h}_{v}^{l}$ is calculated as follows:
@@
where $\mathbf{z}$ is the graph-level representation of $G$ and $\mathbf{\theta_{e}}$ denotes the parameter of our GNN-based encoder. $L$ denotes the number of the graph convolutional layers. The readout function can be implemented in different ways, such as summarizing all nodes’ representations (xu2018powerful, ; yin2022deal, ; yin2023dream, ) or using a virtual node (li2015gated, ; ju2024survey, ; yin2024continuous, ).
@@ 4.1. Second-order Graph ODE
To model high-order correlations in long-term temporal trends,  (rusch2022graph, ) first propose the second-order graph ODE, which is represented as:
@@
where $\left(\mathbf{F}_{\theta}(\mathbf{X},t)\right)_{i}=\mathbf{F}_{\theta}\left(\mathbf{X}_{i}(t),\mathbf{X}_{j}(t),t\right)$ is a learnable coupling function with parameters $\theta$. $\sigma$ is the mapping function, which is implemented with multiplayer perceptions. Due to the unavailability of an analytical solution for Eq. 5, GraphCON (rusch2022graph, ) addresses it through an iterative numerical solver employing a suitable time discretization method. GraphCON utilizes the IMEX (implicit-explicit) time-stepping scheme, an extension of the symplectic Euler method (book) that accommodates systems with an additional damping term.
@@ 4.2. DuSEGO
where $\mathbf{m}_{ij}=\phi_{e}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},||\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}||^{2},a_{ij})$ and $\textbf{m}_{i}=\sum_{j\neq i}\textbf{m}_{ij}$. $\phi_{e}$ and $\phi_{h}$ are the edge and node operations respectively which are commonly approximated by Multilayer Perceptrons (MLPs). Nodes $j$ are the neighbors of node $i$.
@@
(11) $\mathbf{F}_{\theta}(\mathbf{X}^{l},\mathbf{H}^{l},t^{l})=\phi_{h}(\mathbf{h}_{i}^{l},\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij},\mathbf{a}_{i}),\;\mathbf{m}_{ij}=\phi_{m}(\mathbf{h}_{i},\mathbf{h}_{j},||\mathbf{x}_{j}-\mathbf{x}_{i}||^{2},\mathbf{a}_{ij}),$
where $||\mathbf{x}_{j}-\mathbf{x}_{i}||^{2}$ is the squared relative distance between two nodes $v_{i}$ and $v_{j}$, $\phi_{m}$ and $\phi_{h}$ are O(3) steerable MLPs, $\mathbf{a}_{ij}$ and $\mathbf{a}_{i}$ are steerable edge and node attributes, $\mathbf{a}_{i}=\frac{1}{|\mathcal{N}(I)|}\sum_{j\in\mathcal{N}(i)}\mathbf{a}_{ij}$. $\mathcal{N}(i)$ denotes the neighbors of node $v_{i}$.
Steady States of DuSEGO. It is straightforward to see that the steady states $\mathbf{X}^{*}$,$\mathbf{H}^{*}$ of the DuSEGO dynamical system with an autonomous function $\mathbf{F}_{\theta}=\mathbf{F}_{\theta}(\mathbf{X})$ are given by $\mathbf{Y}^{*}\equiv 0$ and $\mathbf{U}^{*}\equiv 0$.
@@
Thus, this interpretation of DuSEGO clearly brings out its relationship with equivariant GNNs. Unlike in standard multi-layer GNNs of the generic that can be thought of as steady states of the underlying ODEs, DuSEGO evolves the underlying node features dynamically in time. Interpreting the multiple GNN layers as iterations at times $t_{n}=n\Delta t$, we observe that the node features in DuSEGO follow the trajectories of the corresponding dynamical system and can explore a richer sampling of the underlying latent feature space, leading to possibly greater expressive power than equivariant GNNs, which might remain in the vicinity of steady states.

### v3

@@ Abstract.
Graph Neural Networks (GNNs) with equivariant properties have achieved significant success in modeling complex dynamic systems and molecular properties. However, their expressiveness ability is limited by: (1) Existing methods often overlook the over-smoothing issue caused by traditional GNN models, as well as the gradient explosion or vanishing problems in deep GNNs. (2) Most models operate on first-order information, neglecting that the real world often consists of second-order systems, which further limits the model’s representation capabilities. To address these issues, we propose the Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO) for equivariant representation. Specifically, DuSEGO applies the dual second-order equivariant graph ordinary differential equations (Graph ODEs) to both graph embeddings and node coordinates simultaneously. Theoretically, we first prove that DuSEGO maintains the equivariant property. Furthermore, we provide theoretical insights showing that DuSEGO effectively alleviates the over-smoothing problem in both feature representation and coordinate update. Additionally, we demonstrate that the proposed DuSEGO mitigates the exploding and vanishing gradients problem, facilitating the training of deep multi-layer GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to baselines.
@@ 1. Introduction
GNNs (kipf2017semisupervised, ) with equivariant properties (satorras2021n, ; han2022equivariant, ; yin2023coco, ; pang2023sa, ) have become essential tools in modeling complex dynamic systems (meng2024deep, ; huang2021equivariant, ; liu2024dynamic, ) and molecular properties (brandstetter2021geometric, ) due to their ability to capture symmetrical relationships within data. Equivariant GNNs are particularly effective in scenarios when maintaining consistency under transformations (e.g., rotations, reflections) is crucial. The consistency of equivariant representations allows these models to accurately reflect the inherent properties of the systems they are designed to study, leading to more reliable and insightful predictions, which is now an essential topic in machine learning (battaglia2016interaction, ; yin2022deal, ; kohler2019equivariant, ).
Despite their successes, existing Equivariant GNNs face notable limitations that constrain their performance and expressiveness. First, the over-smoothing problem (li2018deeper, ; keriven2022not, ; min2020scattering, ; zhang2023investigating, ). The classical Equivariant GNNs (such as EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, )) typically utilize the GNNs (kipf2017semisupervised, ; yin2022dynamic, ) as a backbone, as well as cooperating coordinates for equivariant representation. However, GNNs usually suffer from the over-smoothing problem, leading to the indistinguishable representation of features and coordinates, which is shown in Fig. 1. Second, gradient explosion or vanishing problems (dasoulas2021lipschitz, ; guo2022orthogonal, ; yin2022generic, ; zhang2022would, ). Motivated by the notion that larger models are usually more expressive (kaplan2020scaling, ; ai2023gcn, ), we tend to explore the expressiveness ability by enlarging the model scale. However, as the network depth increases, gradients can either become excessively large or diminish to near zero, which limits the scalability and depth of GNNs. Third, real-world simulation (norcliffe2020second, ). Most Equivariant GNNs (shou2023adversarial, ; thomas2018tensor, ; tang2024merging, ; yinsport, ) primarily focus on first-order information. However, many physical systems are governed by second-order laws (norcliffe2020second, ). The first-order methods overlook the higher-order interactions that are critical in real-world systems, and limit the representational ability. Even though some works (huang2021equivariant, ; liu2024segno, ) try to introduce second-order information into equivariant learning, they simply focus on simulating the position in a dynamic physical system, ignoring the expressive representation for molecular properties prediction.
@@
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node coordinates.
Figure 1. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node coordinates $\mathbf{X}^{n}$ propagated through a EGNN (satorras2021n, ), SEGNN (brandstetter2021geometric, ) and GMN (huang2021equivariant, ) on a N-body system dataset, where we give the definition of the Dirichlet energy in Eq. 4.
To address these challenges, we take an insight into the second-order method in Equivariant GNNs and a framework named Dual Second-order Equivariant Graph Ordinary Differential Equation (DuSEGO). The proposed DuSEGO is a universal framework that transfers the first-order Equivariant GNNs to second-order, which aims to enhance Equivariant GNNs’ expressiveness and address the issues of over-smoothing and gradient stability. First, we prove that the proposed DuSEGO maintains the equivariant properties. Then, we analyze the properties of DuSEGO to reveal how the proposed method solves the over-smoothing, gradient exploding, and vanishing problems, facilitating the training of deep GNNs. Extensive experiments on benchmark datasets validate the superiority of the proposed DuSEGO compared to existing baselines.
The main contributions can be summarized as threefold:
@@ 2. Related Work
Equivariant Graph Neural Networks. Early graph learning methods explored spectral and hypergraph-based techniques that emphasize global topology and higher-order relationships, particularly for tasks such as semi-supervised classification and structure-aware clustering (yin2024continuous, ; ju2024survey, ; wang2025nested, ). Building upon these insights, GNNs (wang2025dynamically, ; yin2025dream, ) with equivariant properties have demonstrated their effectiveness in various tasks, especially for 3D molecular graphs (xian2025molrag, ) and physical symmetry of systems (zhang2023universal, ). The Tensor Field Networks (TFNs) (thomas2018tensor, ) method was first proposed for 3D point clouds, which only leverage the positional information of nodes without any additional features. SE(3)-transformers (fuchs2020se, ) deploy an attention mechanism into TFNs to improve the model performance. EGNN (satorras2021n, ), a simple but effective form of equivariant GNNs, was proposed and has been the backbone of the following research. SEGNN (brandstetter2021geometric, ) incorporates geometric and physical properties into the message-passing framework, significantly enhancing the accuracy and predictive capabilities of EGNN in understanding three-dimensional molecular structures. However, these methods focus on first-order velocity information, even though second-order motion laws govern many physical systems. GMN (huang2021equivariant, ) first introduced the second-order information into equivariant learning, and SEGNO (liu2024segno, ) further applied the second-order ODE for solving the physical dynamic system. However, these methods only target the description of positional state information in the physical world and cannot be applied to learning molecular feature representations. The proposed DuSEGO utilizes the second-order ODE to evaluate the evolution of equivariant representation and position, making it suitable for applications in dynamic physical systems and molecular properties prediction.
@@
Graph Ordinary Differential Equations. Graph Ordinary Differential Equations (Graph ODEs) (poli2019graph, ; ijcai2025p396, ; yin2023messages, ) draw inspiration from Neural Ordinary Differential Equations (chen2018neural, ) and provide a sophisticated mechanism for modeling dynamic changes within graph structures, showing great potential in graph learning (yin2023dream, ; ijcai2025p219, ; sanchez2019hamiltonian, ; yin2023omg, ). A series of works proposes to deploy first-order or second-order ODEs into GNNs to enhance their expressiveness. LG-ODE (huang2020learning, ) learns the dynamics of continuous systems by using models that can handle irregular sampling and incomplete data, improving prediction accuracy and robustness. CG-ODE (huang2021coupled, ) uses coupled ordinary differential equations on graphs to learn and model the dynamics of interacting systems effectively. HOPE (luo2023hope, ) leverages second-order differential equations to capture complex interactions and dynamic behaviors within graph structures. GraphCon (rusch2022graph, ) proposes a novel framework that integrates the dynamics of coupled oscillators with graph theory, analyzing stability across complex networked systems. However, the above works do not consider the equivariance inherent in the graph structure, which is challenging to extend to 3D molecular graphs and the physical symmetries of systems.
@@ 3. Preliminary
Equivariance Let $T_{g}:X\rightarrow X$ be a set of transformations on $X$ for the abstract group $g\in G$. We say a function $\phi:X\rightarrow Y$ is equivariant to $g$ if there exists an equivalent transformation on its output space $S_{g}:Y\rightarrow Y$ such that: $\phi(T_{g}(\mathbf{x}))=S_{g}(\phi(\mathbf{x}))$. The equivariance properties include translations, rotations, and permutations, which ensure that the model’s predictions are consistent regardless of the orientation or position of objects within the input data.
@@
Graph Nerual Networks Given graph $G=(\mathcal{V},\mathcal{E})$, let $\mathbf{h}_{v}^{(k)}$ denote embedding vector for node $v$ at the $l$-th layer. For node $v\in\mathcal{V}$, we aggregate the embeddings from the neighbors of $v$ at layer l1l-1. Then, the node embedding $\mathbf{h}_{v}^{(k)}$ is iteratively updated by combining the embedding of $v$ in the previous layer with the aggregated neighbor embedding. In formulation, $\mathbf{h}_{v}^{l}$ is calculated as follows:
@@
where $\mathbf{z}$ is the graph-level representation of $G$ and $\mathbf{\theta_{e}}$ denotes the parameter of our GNN-based encoder. $L$ denotes the number of the graph convolutional layers. The readout function can be implemented in different ways, such as summarizing all nodes’ representations (xu2018powerful, ; wang2025nested, ) or using a virtual node (li2015gated, ).
@@ 4.1. Second-order Graph ODE
To model high-order correlations in long-term temporal trends,  (rusch2022graph, ) first proposed the second-order graph ODE, which is represented as:
@@
where $\left(\mathbf{F}_{\theta}(\mathbf{X},t)\right)_{i}=\mathbf{F}_{\theta}\left(\mathbf{X}_{i}(t),\mathbf{X}_{j}(t),t\right)$ is a learnable coupling function with parameters $\theta$. $\sigma$ is the mapping function, which is implemented with multiplayer perceptions. Due to the unavailability of an analytical solution for Eq. 5, GraphCON (rusch2022graph, ) addresses it through an iterative numerical solver employing a suitable time discretization method. GraphCON utilizes the IMEX (implicit-explicit) time-stepping scheme, an extension of the symplectic Euler method (wanner1996solving, ) that accommodates systems with an additional damping term.
@@ 4.2. DuSEGO
[…]
@@
[…]
where $\mathbf{x}_{j}-\mathbf{x}_{i}||^{2}$ is the squared relative distance between two nodes $v_{i}$ and $v_{j}$, $\phi_{m}$ and $\phi_{h}$ are O(3) steerable MLPs, $\mathbf{a}_{ij}$ and $\mathbf{a}_{i}$ are steerable edge and node attributes, $\mathbf{a}_{i}=\frac{1}{|\mathcal{N}(I)|}\sum_{j\in\mathcal{N}(i)}\mathbf{a}_{ij}$. $\mathcal{N}(i)$ denotes the neighbors of node $v_{i}$.
Steady States of DuSEGO. It is straightforward to see that the steady states $\mathbf{X}^{*}$$\mathbf{H}^{*}$ of the DuSEGO dynamical system with an autonomous function $\mathbf{F}_{\theta}=\mathbf{F}_{\theta}(\mathbf{X})$ are given by $\mathbf{Y}^{*}\equiv 0$ and $\mathbf{U}^{*}\equiv 0$.
@@
Thus, this interpretation of DuSEGO clearly brings out its relationship with equivariant GNNs. Unlike standard multi-layer GNNs of the generic type, which can be thought of as steady states of the underlying ODEs, DuSEGO dynamically evolves the underlying node features over time. Interpreting the multiple GNN layers as iterations at times $t_{n}=n\Delta t$, we observe that the node features in DuSEGO follow the trajectories of the corresponding dynamical system and can explore a richer sampling of the underlying latent feature space, leading to possibly greater expressive power than equivariant GNNs, which might remain in the vicinity of steady states.
Mechanism of Dynamic Evolution in DuSEGO. DuSEGO evolves node features according to a second-order ordinary differential equation, as defined in Eq. 7. At each iteration, the feature velocity $U_{n}$ is first updated in Eq. 11 based on the previous velocity, the nonlinear transformation of aggregated message passing, and damping terms. The node feature $H_{n}$ is then updated in Eq. 9 by integrating this velocity. This mechanism allows node features to incorporate both instantaneous message-passing effects and accumulated momentum from previous steps. Consequently, feature evolution in DuSEGO is determined not only by current neighborhood information, but also by the historical trajectory of feature changes, enabling a smooth and continuous adaptation of node representations throughout the learning process. The coordinates are updated in the same manner, following an analogous second-order dynamic process as described for the node features.
Intuitive Understanding for Over-smoothing Mitigation of DuSEGO. Different first-order GNNs, where repeated aggregation causes node features to become increasingly similar, resulting in over-smoothing, our second-order ODE framework introduces momentum-like effects into the evolution of both features and coordinates. The second-order terms ensure that updates to each node’s feature and coordinate depend not only on their current state and those of their neighbors, but also on the direction and rate of change from previous steps, analogous to acceleration in physical systems (norcliffe2020second, ; rusch2022graph, ). This added momentum helps preserve the diversity and inertia of node representations and coordinates, preventing rapid or uniform convergence. As a result, the network can be made much deeper without suffering from over-smoothing, as both the features and their geometric arrangement in latent space retain discriminative power across many layers.
4.3. Time Complexity Analysis
The computational complexity of DuSEGO is determined by the coupled evolution of both node representations (features) and node positions (coordinates) via second-order ODEs. At each ODE solver step, the framework performs two independent forward passes through the underlying equivariant GNN backbone—one for the feature branch and one for the coordinate branch.
Let $N$ denote the number of nodes, $d$ the average node degree, $d_{f}$ the hidden feature dimension, and $T$ the number of ODE solver steps. The overall time complexity per iteration depends on the choice of backbone:
  • DuSEGO-EGNN: For the EGNN (satorras2021n, ) backbone, each message passing layer involves lightweight computations such as MLPs and pairwise distance calculations. The complexity of a single forward pass is $\mathcal{O}(N\cdot d\cdot d_{f})$. Since both node features and coordinates are updated independently, the total complexity for $T$ solver steps is:
    $\mathcal{O}(2T\cdot N\cdot d\cdot d_{f}).$
  • DuSEGO-SEGNN: The SEGNN (brandstetter2021geometric, ) backbone incorporates O(3)-steerable message passing, which requires higher-order tensor products and equivariant nonlinearities. The computational cost per forward pass is approximately $\mathcal{O}(N\cdot d\cdot d_{f}^{3})$. Thus, the total complexity for both branches is:
    $\mathcal{O}(2T\cdot N\cdot d\cdot d_{f}^{3}).$